Ed Moura has $33000 invested in stocks paying 5%. How much additional money should he invest in certificates of deposit paying 2% so that the average return on the two investments is 3%?
He should invest $__ in certificates of deposit.

Answer by mananth(16946)

d Moura has $33000 invested in stocks paying 5%.
additional money in certificates of deposit paying 2% =x
so that the average return on the two investments is 3%?

33000*5% +2%x = 3%(33000+x)
multiply by 100
165000 +2x=99000+3x
66000 = 3x-2x
x= $66000 in certificates of deposit paying 2%